# FFHFlow: Diverse and Uncertainty-Aware Dexterous Grasp Generation via Flow Variational Inference

### v3

@@ Abstract
Synthesizing diverse, uncertainty-aware grasps for multi-fingered hands from partial observations remains a critical challenge in robot learning. Prior generative methods struggle to model the intricate grasp distribution of dexterous hands and often fail to reason about shape uncertainty inherent in partial point clouds, leading to unreliable or overly conservative grasps. We propose FFHFlow, a flow-based variational framework that generates diverse, robust multi-finger grasps while explicitly quantifying perceptual uncertainty in the partial point clouds. Our approach leverages a normalizing flow-based deep latent variable model to learn a hierarchical grasp manifold, overcoming the mode collapse and rigid prior limitations of conditional Variational Autoencoders (cVAEs). By exploiting the invertibility and exact likelihoods of flows, FFHFlow introspects shape uncertainty in partial observations and identifies novel object structures, enabling risk-aware grasp synthesis. To further enhance reliability, we integrate a discriminative grasp evaluator with the flow likelihoods, formulating an uncertainty-aware ranking strategy that prioritizes grasps robust to shape ambiguity. Extensive experiments in simulation and real-world setups demonstrate that FFHFlow outperforms state-of-the-art baselines (including diffusion models) in grasp diversity and success rate, while achieving run-time efficient sampling. We also showcase its practical value in cluttered and confined environments, where diversity-driven sampling excels by mitigating collisions (Project Page).
@@ 2 Related Work
Learning-based Grasp Synthesis. Deep generative models such as cVAE[1, 12, 2, 13], auto-regressive models [14], cNF[15, 16, 17], diffusion models [8, 18, 19] and Generative Adversarial Networks (GANs) [20, 21] are widely adopted in grasp synthesis with two-jaw gripper and multi-fingered hands. Among them, some [2, 14] address the partial observation challenge by employing time-intensive shape completion modules, which significantly slow down inference, reducing their practical applicability. Other approaches either assume the availability of complete object point clouds [16], a limitation in real-world scenarios, or only focus on a 2-jaw gripper [17]. Similarly, diffusion models [8, 18, 19] face runtime inefficiencies due to their iterative denoising processes, even though there are some techniques [22, 23] to accelerate it. In contrast, our method, similar to [15, 12], eliminates the need for resource-intensive shape completion models by encoding partial observation information directly into latent variables. This lightweight approach allows our model to effectively handle partial observations while maintaining efficiency and diversity in grasp synthesis. A diverse grasp synthesis enables real-world applications such as grasping in confined space, in clutter and dynamic grasping [24].
@@ Methods
Data: We use only simulated data generated based on a heuristic grasp planner for training in Gazebo Simulator [12]. For training, we use 77 graspable objects filtered from KIT [57] datasets based on their graspability and object type. For testing, we select objects of two levels of difficulty from the KIT and YCB dataset [58] in simulation (Figure 4(a)): (1) Similar: 12 objects from KIT dataset with similar shapes to training objects, serving as ID objects. (2) Novel: 9 objects from YCB dataset with shapes distinct from training objects, often more difficult to grasp, serving as OOD objects. For real-world evaluation, we use 12 unknown objects from YCB dataset. (More detail about the training data and implementation in Section 2 of the appendix.)

### v4

@@ Abstract
Synthesizing diverse, uncertainty-aware grasps for multi-fingered hands from partial observations remains a critical challenge in robot learning. Prior generative methods struggle to model the intricate grasp distribution of dexterous hands and often fail to reason about shape uncertainty inherent in partial point clouds, leading to unreliable or overly conservative grasps. We propose FFHFlow, a flow-based variational framework that generates diverse multi-finger grasps while explicitly quantifying perceptual uncertainty in the partial point clouds. Our approach leverages a normalizing flow-based deep latent variable model to learn a hierarchical grasp manifold, overcoming the mode collapse and rigid prior limitations of conditional Variational Autoencoders (cVAEs). By exploiting the invertibility and exact likelihoods of flows, FFHFlow introspects shape uncertainty in partial observations and identifies novel object structures, enabling risk-aware grasp synthesis. To further enhance reliability, we integrate a discriminative grasp evaluator with the flow likelihoods, formulating an uncertainty-aware ranking strategy that prioritizes grasps robust to shape ambiguity. Extensive experiments in simulation and real-world setups demonstrate that FFHFlow outperforms state-of-the-art baselines (including diffusion models) in grasp diversity and success rate, while achieving run-time efficient sampling. We also showcase its practical value in cluttered and confined environments, where diversity-driven sampling excels by mitigating collisions. (Project Page: https://sites.google.com/view/ffhflow/home/)
@@ 2 Related Work
Learning-based Grasp Synthesis. Deep generative models such as cVAE[1, 12, 2, 13], auto-regressive models [14], cNF[15, 16, 17], diffusion models [8, 18, 19] and Generative Adversarial Networks (GANs) [20, 21] are widely adopted in grasp synthesis with two-jaw gripper and multi-fingered hands. Among them, some [2, 14] address the partial observation challenge by employing time-intensive shape completion modules, which significantly slow down inference, reducing their practical applicability. Other approaches either assume the availability of complete object point clouds [16], a limitation in real-world scenarios, or only focus on a 2-jaw gripper [17]. Similarly, diffusion models [8, 18, 19] face runtime inefficiency due to their iterative denoising processes, even though there are some techniques [22, 23] to accelerate it. In contrast, our method, similar to [15, 12], eliminates the need for resource-intensive shape completion models by encoding partial observation information directly into latent variables. This lightweight approach allows our model to effectively handle partial observations while maintaining efficiency and diversity in grasp synthesis, enabling real-world applications such as grasping in clutter and dynamic grasping [24].
@@ Methods
Data: We use only simulated data generated based on a heuristic grasp planner for training in Gazebo Simulator [12]. For training, we use 77 graspable objects filtered from KIT [57] datasets based on their graspability and object type. For testing, we select objects of two levels of difficulty from the KIT and YCB dataset [58] in simulation (Figure 4(a)): (1) Similar: 12 objects from KIT dataset with similar shapes to training objects, serving as ID objects. (2) Novel: 9 objects from YCB dataset with shapes distinct from training objects, often more difficult to grasp, serving as OOD objects. For real-world evaluation, we use 12 unknown objects from YCB dataset. (More detail in the appendix.)